# L4Q: Parameter Efficient Quantization-Aware Fine-Tuning on Large Language Models

### v4

@@ Abstract
Due to the high memory and computational costs associated with large language models (LLMs), model compression techniques such as quantization, which reduces inference costs, and parameter-efficient fine-tuning (PEFT) methods like Low-Rank Adaptation (LoRA), which reduce training costs, have gained significant popularity. This trend has spurred active research into quantization-aware PEFT techniques, aimed at maintaining model accuracy while minimizing memory overhead during both inference and training. Previous quantization-aware PEFT methods typically follow a two-step approach: first, applying post-training quantization (PTQ) to model weights, followed by PEFT on the quantized model. However, recovering from the quantization error introduced by PTQ through fine-tuning has proven challenging. Additionally, most PTQ-based PEFT methods result in a mixture of low-precision quantized weights and high-precision adapter weights, limiting the efficiency of full quantization during inference. While a previous method attempted to address these issues, it still suffers from limited adaptability due to the constrained LoRA parameter structure required to produce fully-quantized models. To overcome these challenges, we propose L4Q, a method that integrates Quantization-Aware Training (QAT) with LoRA to effectively reduce quantization error. By employing a memory-optimized layer design, L4Q significantly reduces QAT’s memory overhead while producing fully-quantized weights, enabling effective adaptation to downstream tasks. Our experiments demonstrate that this combined approach to quantization and fine-tuning achieves superior accuracy compared to decoupled fine-tuning schemes, particularly in sub-4-bit quantization, positioning L4Q as an efficient QAT solution. Using the LLaMA model families and instructional datasets, we showcase L4Q’s capabilities in language tasks and few-shot learning.
@@ 1 Introduction
Given their impressive scalability, Large Language Models (LLMs) such as GPT, OPT, PaLM, and LLaMA (Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2024; Touvron et al., 2023a; b) are popular in natural language processing. However, their substantial memory and computational demands pose challenges for practical deployment, making model compression (Han et al., 2015) crucial for LLM deployment. Quantization is a prominent method that reduces model size by lowering the bit precision of model parameters (Hubara et al., 2017), so LLM quantization has been actively studied (Liu et al., 2024; Xiao et al., 2023; Frantar et al., 2023; Dettmers & Zettlemoyer, 2023). These quantization methods are generally divided into two categories: post-training quantization (PTQ) and quantization-aware training (QAT). QAT effectively reduces quantization error by integrating quantization into the training process, where both model weights and quantization parameters, are trained together (Esser et al., 2020; Bhalgat et al., 2020). However, applying QAT to LLM quantization is challenging due to the significant memory overhead it incurs during training. As a result, PTQ, which applies quantization without retraining the pre-trained model weights and with minimal calibration of the quantization parameters, is widely adopted for LLM quantization (Xiao et al., 2023; Lin et al., 2024; Heo et al., 2024).
Concurrently, to enhance the problem-solving abilities of LLMs for specific applications, fine-tuning pre-trained LLMs on downstream tasks is crucial as it improves accuracy on target tasks and related tasks (Wei et al., 2022; Scialom et al., 2022). However, fine-tuning is a resource-intensive process due to the large number of model weight parameters involved. Parameter-efficient fine-tuning (PEFT) addresses this issue (Hu et al., 2022; Li & Liang, 2021; Liu et al., 2022a; Wang et al., 2023) by training a small subset of parameters while freezing the majority of pre-trained weights. One of the most prominent techniques within PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which inserts trainable rank decomposition matrices into each layer to represent updates to the frozen weights.
@@
Figure 1: A diagram of model fine-tuning and quantization represented with an example of LLaMA-1 7B model. L4Q produces fast and accurate quantized model.
@@
The integration of quantization and PEFT holds significant potential for developing highly efficient and accurate LLMs for downstream tasks. Recent research has introduced quantization-aware PEFT approaches to achieve high-quality quantized models (Dettmers et al., 2024b; Kim et al., 2024; Xu et al., 2023; Li et al., 2024). Previous works involve a two-stage optimization strategy: first, a PTQ technique, such as GPTQ (Frantar et al., 2023), is applied to pre-trained LLMs for compression. Then, these quantized LLMs undergo PEFT, such as LoRA, where quantized weights are kept fixed and only the LoRA parameters are fine-tuned. While fine-tuning can mitigate the effects of quantization errors, separating quantization and fine-tuning into distinct state hinders the models from achieving the best accuracy. Furthermore, as high-precision LoRA parameters are adopted alongside the quantized weight matrix, these methods eventually produce mixed-precision models, which limits the efficiency of full quantization during inference. Recently, QA-LoRA (Xu et al., 2023) addresses this issue by strictly constraining the LoRA parameter structure to integrate with quantization parameters, but this constraint can limit the fine-tuning capability.
@@
In this paper, we propose a novel quantization-aware fine-tuning technique, named L4Q (Low-rank adaptive Learning quantization for LLMs). L4Q addresses the limitations of PTQ-based PEFT methods by introducing QAT alongside LoRA. While QAT and LoRA have advantages in reducing quantization error and enabling memory-efficient training, respectively, their straightforward integration diminishes the benefits of each approach. Therefore, L4Q carefully integrates these two approaches to properly leverage their advantages. First, L4Q applies the quantization process after fully combining the model weights and LoRA parameters in the linear layer. This approach produces a fully-quantized model that enables memory-efficient and fast inference without limiting the training capabilities of either QAT or LoRA. Moreover, to preserve the memory-efficient nature of LoRA during training, we design the backpropagation path of L4Q to eliminate the need to store weight gradients required for QAT. Finally, the full integration of QAT and LoRA in the proposed L4Q allows for the joint optimization of both the quantization and LoRA parameters, thereby improving of quantized LLMs. As a result, L4Q significantly improves the accuracy of quantized models while maintaining low memory costs during both inference and training, and achieves inference speed comparable to state-of-the-art approaches, making it a more effective solution compared to previous works, as illustrated in Figure 1.
@@ 2.2 Quantization
Uniform quantization is a widely used quantization scheme due to its simplicity and broad compatibility with various computing kernels and hardware units (Liu et al., 2022b). Therefore, considering the adaptability on the broad quantization scheme, including both weight-only and activation quantization scenarios, we refer to the term ‘quantization’ specifically as uniform quantization throughout this paper. A common practice is to organize a quantization group consisting of a certain number of consecutive weight elements that share the same quantization scale $s$ and bias $b$, which control the quantization step size and zero point, respectively.
@@
$\tilde{w}=R(clamp(\frac{W-b}{s},Q_{n},Q_{p})),\quad W_{q}=\tilde{w}\times s+b$ (3)
Here, $\tilde{w}$ denotes the quantized integer value obtained through a sequence of division, clamping, and rounding. Clamping is applied within the range $Q_{n}=-2^{n-1}$ to $Q_{p}=2^{n-1}-1$, where $n$ is the bit-width, followed by the rounding function $R$. $W_{q}$ represents the dequantized version of the quantized weight, which is adjusted using $s$ and $b$ from $\tilde{w}$ to approximate the original weight.
During QAT, the straight-through estimator (STE) approximates the derivative of the rounding function with an identity function (Bengio et al., 2013; Choi et al., 2018; Esser et al., 2020), enabling gradients to propagate through non-differentiable rounding operations and allowing effective weight parameter training. Building on conventional QAT, LSQ (Esser et al., 2020) and LSQ+ (Bhalgat et al., 2020) extend this process by training quantization parameters $s$ and $b$, alongside model weights. This tuning scheme provides finer control over quantization, improving model accuracy. The quantization parameters tuning during backpropagation proceeds as follows, where $w$ denotes the value $\frac{W-b}{s}$:
$\frac{\partial L}{\partial s}=\frac{\partial W_{q}}{\partial s}\frac{\partial L}{\partial W_{q}}\enspace s.t.\enspace\frac{\partial W_{q}}{\partial s}=-w+\tilde{w},\quad\mathrm{if}\enspace Q_{n}\leq w\leq Q_{p}$ (4)
$\frac{\partial L}{\partial b}=\frac{\partial W_{q}}{\partial b}\frac{\partial L}{\partial W_{q}}\enspace s.t.\enspace\frac{\partial W_{q}}{\partial b}=1,\quad\quad\;\mathrm{if}\enspace w<Q_{n}\enspace or\enspace w>Q_{p}$ (5)
@@ 2.3 LLM Quantization
Quantization compress LLMs by lowering the bit precision of model parameters (Hubara et al., 2017). However, a key challenge with quantization is the introduction of errors that reduce model accuracy, leading to extensive research aimed at mitigating these losses through calibration or training. A notable examples of PTQ for LLM compression are GPTQ(Frantar et al., 2023) and OmniQuant (Shao et al., 2023). In contrast, QAT integrates quantization into the training process, adaptively training model parameters to account for quantization errors during training, ensuring that the quantized model retains much of its accuracy and functionality through training.
Despite its advantages, QAT faces challenges, primarily due to its high training overhead, which limits its use in resource-intensive models like LLMs. The memory overhead of QAT stems from storing weight gradients and their optimizer states, each requiring multiple times the memory of the weights. For example, applying QAT to a 7B model that fine-tunes all weight parameters requires at least 14GB for the model weights, 14GB for the gradients, 28GB for optimizer states, and additional dozens of GB for activations. This often exceeds the memory capacity of a single GPU, which typically offers a maximum of 80GB of memory. Therefore, previous works have primarily focused on the application of PTQ for LLM compression, while research on QAT remains in its early stages, especially in terms of improving training efficiency.
@@ 2.4 Quantization-Aware PEFT
To improve the accuracy of quantized LLMs, recent research has introduced quantization-aware PEFT approaches (Dettmers et al., 2024b; Kim et al., 2024; Xu et al., 2023; Li et al., 2024). Among these, QLoRA (Dettmers et al., 2024b), QA-LoRA (Xu et al., 2023), and LoftQ (Li et al., 2024) stand out as notable methods. As illustrated in Figure 2, QLoRA begins by applying PTQ to a pre-trained model. After this initial quantization, LoRA fine-tuning is performed, with the quantized weight parameters kept frozen. The LoRA parameters are fine-tuned using higher precision formats such as bfloat16 or float16, allowing the method to correct quantization errors during the fine-tuning. However, QLoRA introduces computational inefficiencies during inference due to the additional forward path on LoRA parameters. This inefficiency arises because the high-precision LoRA parameters and low-precision quantized weights cannot be merged into low-precision values. We further examine the impact of this unmerged LoRA path on inference efficiency by comparing the speed of fully-quantized models with mixed-precision models in Section 4.
QA-LoRA (Xu et al., 2023) addresses the issue of high-precision LoRA parameters by modifying the structure of the LoRA matrix, allowing these parameters to be integrated with the quantized weights after training (Figure 2). The input dimension of the LoRA matrix $A$ is set to the number of weight quantization groups. This adjustment ensures that each element of $BA$ corresponds directly with individual quantization groups, enabling the LoRA parameters to be seamlessly integrated into the quantization bias as $b^{\prime}=b-\alpha BA$ at the end of training. However, this solution presents a new challenge: the constrained LoRA structure in this setup limits the model’s ability to achieve optimal accuracy during the PEFT stage.
A broader issue with previous quantization-aware PEFT approaches is that quantization and fine-tuning are performed sequentially rather than simultaneously. Starting the fine-tuning process from a pre-quantized model with inherent quantization errors is less optimal than starting from a pre-trained model. While LoftQ attempts to address these quantization errors by approximating them using LoRA with iterative singular-value decomposition (SVD), the limitations on subsequent adaptation persist, making it challenging to fully recover and optimize model performance during fine-tuning. These challenges highlight the need for continued research to improve quantization-aware PEFT techniques, aiming to enhance both the quantization and PEFT processes for better accuracy and efficiency in LLMs.
@@ 3 Methods
3.1 Straightforward Integration of QAT and LoRA
QAT-LoRA  One of the key principles in our proposed L4Q scheme is the integration of the QAT and LoRA to facilitate the simultaneous calibration for quantization and fine-tuning on downstream tasks. To achieve this, we begin with a straightforward integration of QAT and LoRA, referred to as QAT-LoRA, which serves as our baseline approach for combining QAT and PEFT. In QAT-LoRA, pre-trained weights are frozen, while LoRA parameters are added to the linear layers (Figure 2). Additionally, quantization scales and bias parameters are introduced, similar to advanced QAT techniques like LSQ, which are crucial for calibrating the quantization function. Freezing the weights reduces the need for optimizer states, while a small number of LoRA and quantization parameters are introduced to approximate updates to the weight matrix and to update the quantization function, respectively. This results in more efficient memory usage compared to standard QAT. Detailed analysis results of the memory efficiency of QAT-LoRA is further discussed in Section 4.
Limitations of QAT-LoRA  However, several issues arise with this straightforward integration of QAT and LoRA, where quantized weights and LoRA parameters remain distinct. First, although freezing the pre-trained weights eliminates the need for optimizer states, weight gradients $\frac{\partial L}{\partial W_{q}}$ must still be stored to update the quantization parameters, as shown in Equations  45. As a result, QAT-LoRA still incurs memory overhead from weight gradients, undermining the memory efficiency benefits of LoRA fine-tuning. Secondly, QAT-LoRA produces a mixed-precision model with both quantized weights and high-precision LoRA parameters. This mixed-precision approach negates the advantages of LLM quantization, similar to previous methods such as QLoRA and LoftQ discussed in Section 2.4. Lastly, the gradient updates for quantization and LoRA parameters are decoupled, limiting the potential for comprehensive optimization across the model. As outlined in Equations  45, updates to the quantization parameters rely on the quantized weight matrix $W_{q}$, while updates to the LoRA parameters depend on weights $A$ and $B$. This limits the effectiveness of model training, as it prevents holistic adjustments where changes in LoRA parameters could directly influence quantization adjustments and vice versa.

### v5

@@ Abstract
Due to the high memory and computational costs associated with large language models (LLMs), model compression techniques such as quantization, which reduces inference costs, and parameter-efficient fine-tuning (PEFT) methods like Low-Rank Adaptation (LoRA), which reduce training costs, have gained significant popularity. This trend has spurred active research into quantization-aware PEFT techniques, aimed at maintaining model accuracy while minimizing memory overhead during both inference and training. Previous quantization-aware PEFT methods typically apply post-training quantization (PTQ) to pre-trained LLMs, followed by PEFT to recover accuracy loss. Meanwhile, this approach has limitations in recovering the accuracy loss. In this paper, we propose L4Q, a method that integrates Quantization-Aware Training (QAT) with LoRA. By employing a memory-optimized layer design, L4Q significantly reduces QAT’s memory overhead, making its training cost comparable to LoRA, while preserving the advantage of QAT in producing fully quantized LLMs with high accuracy. Our experiments demonstrate that this combined approach to quantization and fine-tuning achieves superior accuracy compared to decoupled fine-tuning schemes, particularly in 4-bit and 3-bit quantization, positioning L4Q as an efficient QAT solution. Using the LLaMA and Mistral models with instructional datasets, we showcase L4Q’s capabilities in language tasks and few-shot learning.
L4Q: Parameter Efficient Quantization-Aware Fine-Tuning
on Large Language Models
Hyesung Jeon Seoul National University hjeon2k@snu.ac.kr                        Yulhwa Kim Sungkyunkwan University yulhwakim@skku.edu                        Jae-Joon Kim Seoul National University kimjaejoon@snu.ac.kr
@@ 1 Introduction
Figure 1: A diagram of accuracy and inference speedup of the quantized or fine-tuned LLaMA-1 7B models. L4Q produces fast and accurate quantized model.
Given their impressive scalability, Large Language Models (LLMs) such as GPT, OPT, PaLM, and LLaMA (Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2024; Touvron et al., 2023a, b) have become popular in natural language processing. However, their substantial memory and computational demands pose challenges for practical deployment, making model compression (Han et al., 2015) crucial for LLM deployment. Quantization is a prominent method that reduces model size by lowering the bit precision of model parameters (Hubara et al., 2017), so LLM quantization has been actively studied (Liu et al., 2024; Xiao et al., 2023; Frantar et al., 2023; Dettmers and Zettlemoyer, 2023). These quantization methods are generally divided into two categories: quantization-aware training (QAT) and post-training quantization (PTQ). QAT effectively reduces the quantization error by integrating quantization into the training process, where both the model weights and the quantization parameters are trained together (Esser et al., 2020; Bhalgat et al., 2020). However, applying QAT to LLM is challenging due to its significant memory overhead. As a result, PTQ, which applies quantization without retraining the entire pre-trained model weights and with minimal calibration of the quantization parameters, is widely adopted for LLM quantization (Xiao et al., 2023; Lin et al., 2024; Heo et al., 2024).
Concurrently, to enhance the problem-solving abilities of LLMs for specific applications, fine-tuning pre-trained LLMs on downstream tasks is crucial as it improves accuracy on target tasks and related tasks (Wei et al., 2022; Scialom et al., 2022). However, fine-tuning is a resource-intensive process due to the large number of model weight parameters involved. Parameter-efficient fine-tuning (PEFT) addresses this issue (Hu et al., 2022; Li and Liang, 2021; Liu et al., 2022a; Wang et al., 2023) by training a small subset of parameters while freezing the majority of pre-trained weights. One of the most prominent techniques within PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which inserts trainable rank decomposition matrices into each layer to represent updates to the frozen weights.
@@
The integration of quantization and PEFT holds significant potential for developing efficient and accurate LLMs for downstream tasks. Recent research has introduced quantization-aware PEFT approaches to achieve high-quality quantized models (Dettmers et al., 2024b; Kim et al., 2024; Xu et al., 2023; Li et al., 2024). Previous works involve a two-stage optimization strategy: first, a PTQ technique, such as GPTQ (Frantar et al., 2023), is applied to pre-trained LLMs for compression. Then, these quantized LLMs undergo PEFT, such as LoRA, where quantized weights are kept fixed and only the LoRA parameters are fine-tuned. While fine-tuning can mitigate the effects of quantization errors, separating quantization and fine-tuning into distinct stages hinders the models from achieving the best accuracy. Furthermore, as high-precision LoRA parameters are adopted alongside the quantized weight matrix, these methods eventually produce mixed-precision models, which limits the efficiency of full quantization during inference. Recently, QA-LoRA (Xu et al., 2023) addresses this issue by strictly constraining the LoRA parameter structure to integrate with quantization parameters, but this constraint can limit the fine-tuning capability.
@@
In this paper, we propose a novel quantization-aware fine-tuning technique, named L4Q (Low-rank adaptive Learning quantization for LLMs). L4Q addresses the limitations of PTQ-based PEFT methods by introducing QAT alongside LoRA. While QAT have advantages in reducing quantization error and LoRA enables memory-efficient training, their straightforward integration diminishes the benefits of each approach. Therefore, L4Q carefully integrates these two approaches to properly leverage their advantages. First, L4Q applies the quantization process after fully combining the model weights and LoRA parameters in the linear layer. This approach produces a fully-quantized model that enables memory-efficient and fast inference without limiting the training capabilities of either QAT or LoRA. Moreover, to preserve the memory-efficient nature of LoRA during training, we design the backpropagation path of L4Q to eliminate the need to store weight gradients required for QAT. Finally, the full integration of QAT and LoRA in the proposed L4Q allows for the joint optimization of both the quantization and LoRA parameters, thereby improving the quality of the quantized LLMs. As a result, L4Q significantly improves the accuracy of quantized models while maintaining low memory costs during both inference and training, and achieves inference speed comparable to state-of-the-art approaches, making it a more effective solution compared to previous works, as illustrated in Figure 1.
@@ 2.2 Quantization
Uniform quantization is a widely used quantization scheme due to its simplicity and broad compatibility with various computing kernels and hardware units (Liu et al., 2022b). Therefore, we refer to the term ‘quantization’ specifically as uniform quantization throughout this paper. A common practice is to organize a quantization group consisting of a certain number of consecutive weight elements that share the same quantization scale $s$ and zero-point (bias) $b$.
@@
$\tilde{w}=\text{round}(\text{clamp}(\frac{W-b}{s},Q_{N},Q_{P}))$ (3)
Here, $\tilde{w}$ denotes the quantized integer value. Clamping is applied within the range $Q_{N}=-2^{n-1}$ to $Q_{P}=2^{n-1}-1$, where $n$ is the bit-width, followed by the rounding function. We also note that $W_{q}:=\tilde{w}\times s+b$ represents the dequantized version of the quantized weight, which is adjusted using $s$ and $b$ from $\tilde{w}$ to approximate the original weight.
During QAT, the straight-through estimator (STE) approximates the derivative of the rounding function with an identity function (Bengio et al., 2013; Choi et al., 2018; Esser et al., 2020), enabling gradients to propagate through non-differentiable rounding operations and allowing effective weight parameter training. LSQ (Esser et al., 2020) and LSQ+ (Bhalgat et al., 2020) extend this process by training quantization parameters $s$ and $b$, alongside the model weights. This tuning scheme provides finer control over quantization, improving model accuracy. The quantization parameters tuning during backpropagation proceeds by using the chain rule via $W_{q}$. This is denoted as follows:
$\frac{\partial L}{\partial s}=\frac{\partial L}{\partial W_{q}}\frac{\partial W_{q}}{\partial s},\quad\frac{\partial L}{\partial b}=\frac{\partial L}{\partial W_{q}}\frac{\partial W_{q}}{\partial b}$ (4)
As a consequence, the backpropagation process requires the weight gradient $\frac{\partial L}{\partial W_{q}}$ and the computation of the terms $\frac{\partial W_{q}}{\partial s},\frac{\partial W_{q}}{\partial b}$ regarding the non-linear STE function.
$\frac{\partial W_{q}}{\partial s}=-w+\tilde{w},\quad\mathrm{if}\enspace Q_{N}\leq w\leq Q_{P}$ (5)
$\frac{\partial W_{q}}{\partial b}=1,\quad\mathrm{if}\enspace w<Q_{N}\enspace or\enspace w>Q_{P}$ (6)
@@ 2.3 LLM Quantization
Quantization compress LLMs by lowering the bit precision of model parameters (Hubara et al., 2017). A key challenge is the introduction of quantization errors that reduce model accuracy, leading to extensive research aimed at mitigating these losses through calibration or training. A notable examples of PTQ for LLM compression are GPTQ (Frantar et al., 2023) and OmniQuant (Shao et al., 2023). In contrast, QAT integrates quantization into the training process, adaptively training model parameters to account for quantization errors during training, ensuring that the quantized model retains much of its accuracy and functionality through training. Despite its advantages, QAT faces challenges, primarily due to its high training overhead, which limits its use in resource-intensive models like LLMs. The memory overhead of QAT stems from storing weight gradients and their optimizer states, each requiring multiple times the memory of the weights. Hence, even applying QAT to a 7B model requires approximately 80GB of memory.
@@ 2.4 Quantization-Aware PEFT
To improve the accuracy of quantized LLMs, recent research has introduced quantization-aware PEFT approaches (Dettmers et al., 2024b; Kim et al., 2024; Xu et al., 2023; Li et al., 2024). Among these, QLoRA (Dettmers et al., 2024b), QA-LoRA (Xu et al., 2023), and LoftQ (Li et al., 2024) stand out as notable methods. As illustrated in Figure 2, QLoRA begins by applying PTQ to a pre-trained model. After this initial quantization, LoRA fine-tuning is performed, with the quantized weight parameters kept frozen. This allows the method to correct quantization errors during the fine-tuning. However, QLoRA introduces computational inefficiencies during inference due to the additional forward path on LoRA parameters. This inefficiency arises because the high-precision LoRA parameters and low-precision quantized weights cannot be merged into low-precision values. Advanced methods (Li et al., 2024; Qin et al., 2024) that build upon QLoRA and share its layer structure also suffer from this issue. We further examine the impact of this unmerged LoRA path on inference efficiency by comparing the speed of fully-quantized models with mixed-precision models in Section 4.
QA-LoRA (Xu et al., 2023) addresses the issue of high-precision LoRA parameters by modifying the structure of the LoRA matrix, allowing these parameters to be integrated with the quantized weights after training (Figure 2). The input dimension of the LoRA matrix $A$ is set to the number of weight quantization groups. This adjustment ensures that each element of $BA$ corresponds directly with individual quantization groups, enabling the LoRA parameters to be seamlessly integrated into the quantization bias as $b^{\prime}=b-\alpha BA$ at the end of training. Hence, QA-LoRA shares the same objective as our work. However, this solution presents a new challenge: the constrained LoRA structure in this setup limits the model’s ability to achieve optimal accuracy during the PEFT stage.
A broader issue with existing quantization-aware PEFT methods is that fine-tuning begins from a pre-quantized model with inherent quantization errors, which is suboptimal compared to starting from a pre-trained model. LoftQ attempts to mitigate these errors by approximating them with LoRA using iterative singular-value decomposition (SVD). However, this approach still cannot achieve a single forward path due to the high-precision LoRA parameters, limiting subsequent adaptation. These challenges underscore the need for further research to improve quantization-aware PEFT techniques, focusing on enhancing both quantization and PEFT processes for better accuracy and inference efficiency in LLMs.
@@ 3 Methods
3.1 Straight Integration of QAT and LoRA
QAT-LoRA  One of the key principles in our proposed L4Q scheme is the integration of the QAT and LoRA to facilitate the simultaneous calibration for quantization and fine-tuning on downstream tasks. To achieve this, we begin with a straightforward integration of QAT and LoRA, referred to as QAT-LoRA, which serves as our baseline approach for combining QAT and PEFT.
In QAT-LoRA, pre-trained weights are frozen, while LoRA parameters are added to the linear layers (Figure 2). Additionally, quantization scales and bias parameters are introduced, similar to advanced QAT techniques like LSQ, which are crucial for calibrating the quantization function. Freezing the weights reduces the need for optimizer states, while a small number of LoRA and quantization parameters are introduced to approximate updates to the weight matrix and to update the quantization function, respectively. This results in more efficient memory usage compared to standard QAT. Detailed analysis results of the memory efficiency of QAT-LoRA is further discussed in Section 4.
Limitations of QAT-LoRA  However, several issues arise with this straightforward integration of QAT and LoRA, where quantized weights and LoRA parameters remain distinct. First, although freezing the pre-trained weights eliminates the need for optimizer states, weight gradients $\frac{\partial L}{\partial W_{q}}$ must still be stored to update the quantization parameters, as shown in Equation 5 and Equation 6. As a result, QAT-LoRA still incurs memory overhead from weight gradients, undermining the memory efficiency benefits of LoRA fine-tuning. Secondly, QAT-LoRA produces a mixed-precision model with both quantized weights and high-precision LoRA parameters. This mixed-precision approach negates the advantages of LLM quantization, similar to previous methods such as QLoRA and LoftQ discussed in Section 2.4. Lastly, the gradient updates for quantization and LoRA parameters are decoupled, limiting the potential for comprehensive optimization across the model. As outlined in Equations  5 and  6, updates to the quantization parameters rely on the quantized weight matrix $W_{q}$, while updates to the LoRA parameters depend on weights $A$ and $B$. This limits the effectiveness of model training, as it prevents holistic adjustments where changes in LoRA parameters could directly influence quantization adjustments and vice versa.